OpenStudy (anonymous):

Troy was trying to catch up with the rest of his group, who had left earlier on a kayaking trip. Troy paddled 8 hours downstream, covering 72 miles before realizing that he had passed his group’s campground. He had to paddle back upstream for 18 miles to get back to his group’s campground. If the trip back upstream took Troy 6 hours, how fast was the river flowing (in mph)? A. 2 mph B. 3 mph C. 6 mph D. 9 mph

OpenStudy (anonymous):

@xapproachesinfinity

OpenStudy (anonymous):

@Nurali

OpenStudy (nurali):

so let b=speed of boat c= speed of surrent b+c= downstream b-c=upstream

OpenStudy (nurali):

set up 2 equations 8(b+c)=72 6(b-c)=18

OpenStudy (anonymous):

Hmm

OpenStudy (nurali):

so divide both sides by 8 in the first equation and divide both sides by 6 in the second equaiton

OpenStudy (nurali):

and multiply one of the equation by -1 then u can use elimination and solve for c

OpenStudy (nurali):

ur answer is 3 mph

OpenStudy (anonymous):

Ohhh!!!